Cl-o bond order in perchlorate ion will be.....Please explain how to calculate bond order for polyatomic ions like aboveAlso tell forCLO4^- PO4^3- SO4^2- SiO4^4-

Grade:12

1 Answers

Arun
6 years ago
  • ClO4- : (Valency of one Peripheral atom Oxygen = 2, Charge on acid radical = -1, Total Number of Peripheral atoms = 04), Therefore B.O. = 2 + (-1/4) = 1.75                        
  • ClO3- : (Valency of one Peripheral atom Oxygen = 2, Charge on acid radical = -1, Total Number of Peripheral atoms = 03), Therefore B.O. = 2 + (-1/3) = 1.66                        
  • ClO2- : (Valency of one Peripheral atom Oxygen = 2, Charge on acid radical = -1, Total Number of Peripheral atoms = 02), Therefore B.O. = 2 + (-1/2) = 1.5